\begin{thebibliography}{26}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Abdulmumin et~al.(2025)Abdulmumin, Agostinelli, Alumäe, Anastasopoulos, Ashwin, Bentivogli, Bojar, Borg, Bougares, Cattoni, Cettolo, Chen, Chen, Dabre, Estève, Federico, Gaido, Javorský, Kasztelnik, Lam, Liu, Matusov, Maurya, McCrae, Mdhaffar, Moslem, Murray, Nakamura, Negri, Niehues, Ojha, Ortega, Papi, Pecina, Polák, Połeć, Savoldi, Sethiya, Sikasote, Sperber, Stüker, Sudoh, Thompson, Turchi, Waibel, Wilken, Zevallos, Zouhar, and Züfle}]{iwslt2025}
Idris Abdulmumin, Victor Agostinelli, Tanel Alumäe, Antonios Anastasopoulos, Ashwin, Luisa Bentivogli, Ondřej Bojar, Claudia Borg, Fethi Bougares, Roldano Cattoni, Mauro Cettolo, Lizhong Chen, William Chen, Raj Dabre, Yannick Estève, Marcello Federico, Marco Gaido, Dávid Javorský, Marek Kasztelnik, and 30 others. 2025.
\newblock Findings of the iwslt 2025 evaluation campaign.
\newblock In \emph{Proceedings of the 22nd International Conference on Spoken Language Translation (IWSLT 2025)}, Vienna, Austia (in-person and online). Association for Computational Linguistics.
\newblock To appear.

\bibitem[{Agarwal et~al.(2023)Agarwal, Agrawal, Anastasopoulos, Bentivogli, Bojar, Borg, Carpuat, Cattoni, Cettolo, Chen, Chen, Choukri, Chronopoulou, Currey, Declerck, Dong, Duh, Est{\`e}ve, Federico, Gahbiche, Haddow, Hsu, Mon~Htut, Inaguma, Javorsk{\'y}, Judge, Kano, Ko, Kumar, Li, Ma, Mathur, Matusov, McNamee, P.~McCrae, Murray, Nadejde, Nakamura, Negri, Nguyen, Niehues, Niu, Kr.~Ojha, E.~Ortega, Pal, Pino, van~der Plas, Pol{\'a}k, Rippeth, Salesky, Shi, Sperber, St{\"u}ker, Sudoh, Tang, Thompson, Tran, Turchi, Waibel, Wang, Watanabe, and Zevallos}]{agrawal-etal-2023-findings}
Milind Agarwal, Sweta Agrawal, Antonios Anastasopoulos, Luisa Bentivogli, Ond{\v{r}}ej Bojar, Claudia Borg, Marine Carpuat, Roldano Cattoni, Mauro Cettolo, Mingda Chen, William Chen, Khalid Choukri, Alexandra Chronopoulou, Anna Currey, Thierry Declerck, Qianqian Dong, Kevin Duh, Yannick Est{\`e}ve, Marcello Federico, and 43 others. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.iwslt-1.1} {{FINDINGS} {OF} {THE} {IWSLT} 2023 {EVALUATION} {CAMPAIGN}}.
\newblock In \emph{Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, pages 1--61, Toronto, Canada (in-person and online). Association for Computational Linguistics.

\bibitem[{Ahmad et~al.(2024)Ahmad, Anastasopoulos, Bojar, Borg, Carpuat, Cattoni, Cettolo, Chen, Dong, Federico, Haddow, Javorsk{\'y}, Krubi{\'n}ski, Lam, Ma, Mathur, Matusov, Maurya, McCrae, Murray, Nakamura, Negri, Niehues, Niu, Ojha, Ortega, Papi, Pol{\'a}k, Posp{\'i}{\v{s}}il, Pecina, Salesky, Sethiya, Sarkar, Shi, Sikasote, Sperber, St{\"u}ker, Sudoh, Thompson, Waibel, Watanabe, Wilken, Zem{\'a}nek, and Zevallos}]{ahmad-etal-2024-findings}
Ibrahim~Said Ahmad, Antonios Anastasopoulos, Ond{\v{r}}ej Bojar, Claudia Borg, Marine Carpuat, Roldano Cattoni, Mauro Cettolo, William Chen, Qianqian Dong, Marcello Federico, Barry Haddow, D{\'a}vid Javorsk{\'y}, Mateusz Krubi{\'n}ski, Tsz~Kin Lam, Xutai Ma, Prashant Mathur, Evgeny Matusov, Chandresh Maurya, John McCrae, and 25 others. 2024.
\newblock \href {https://doi.org/10.18653/v1/2024.iwslt-1.1} {{FINDINGS} {OF} {THE} {IWSLT} 2024 {EVALUATION} {CAMPAIGN}}.
\newblock In \emph{Proceedings of the 21st International Conference on Spoken Language Translation (IWSLT 2024)}, pages 1--11, Bangkok, Thailand (in-person and online). Association for Computational Linguistics.

\bibitem[{Bentivogli et~al.(2021)Bentivogli, Cettolo, Gaido, Karakanta, Martinelli, Negri, and Turchi}]{bentivogli2021cascade}
Luisa Bentivogli, Mauro Cettolo, Marco Gaido, Alina Karakanta, Alberto Martinelli, Matteo Negri, and Marco Turchi. 2021.
\newblock Cascade versus direct speech translation: Do the differences still make a difference?
\newblock \emph{arXiv preprint arXiv:2106.01045}.

\bibitem[{B{\'e}rard et~al.(2016)B{\'e}rard, Pietquin, Servan, and Besacier}]{berard2016listen}
Alexandre B{\'e}rard, Olivier Pietquin, Christophe Servan, and Laurent Besacier. 2016.
\newblock Listen and translate: A proof of concept for end-to-end speech-to-text translation.
\newblock \emph{arXiv preprint arXiv:1612.01744}.

\bibitem[{Communication et~al.(2023)Communication, Barrault, Chung, Meglioli, Dale, Dong, Duquenne, Elsahar, Gong, Heffernan, Hoffman, Klaiber, Li, Licht, Maillard, Rakotoarison, Sadagopan, Wenzek, Ye, Akula, Chen, Hachem, Ellis, Gonzalez, Haaheim, Hansanti, Howes, Huang, Hwang, Inaguma, Jain, Kalbassi, Kallet, Kulikov, Lam, Li, Ma, Mavlyutov, Peloquin, Ramadan, Ramakrishnan, Sun, Tran, Tran, Tufanov, Vogeti, Wood, Yang, Yu, Andrews, Balioglu, Costa-jussà, Celebi, Elbayad, Gao, Guzmán, Kao, Lee, Mourachko, Pino, Popuri, Ropers, Saleem, Schwenk, Tomasello, Wang, Wang, and Wang}]{barrault2023seamlessm4t}
Seamless Communication, Loïc Barrault, Yu-An Chung, Mariano~Cora Meglioli, David Dale, Ning Dong, Paul-Ambroise Duquenne, Hady Elsahar, Hongyu Gong, Kevin Heffernan, John Hoffman, Christopher Klaiber, Pengwei Li, Daniel Licht, Jean Maillard, Alice Rakotoarison, Kaushik~Ram Sadagopan, Guillaume Wenzek, Ethan Ye, and 49 others. 2023.
\newblock Seamlessm4t: Massively multilingual \& multimodal machine translation.
\newblock \emph{arXiv preprint arXiv:2308.11596}.

\bibitem[{E.~Ortega et~al.(2023)E.~Ortega, Zevallos, and Chen}]{e-ortega-etal-2023-quespa}
John E.~Ortega, Rodolfo Zevallos, and William Chen. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.iwslt-1.23} {{QUESPA} submission for the {IWSLT} 2023 dialect and low-resource speech translation tasks}.
\newblock In \emph{Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, pages 261--268, Toronto, Canada (in-person and online). Association for Computational Linguistics.

\bibitem[{Gow-Smith et~al.(2023)Gow-Smith, Berard, Zanon~Boito, and Calapodescu}]{gow-smith-etal-2023-naver}
Edward Gow-Smith, Alexandre Berard, Marcely Zanon~Boito, and Ioan Calapodescu. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.iwslt-1.10} {{NAVER} {LABS} {E}urope{'}s multilingual speech translation systems for the {IWSLT} 2023 low-resource track}.
\newblock In \emph{Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, pages 144--158, Toronto, Canada (in-person and online). Association for Computational Linguistics.

\bibitem[{Jia et~al.(2019)Jia, Weiss, Biadsy, Macherey, Johnson, Chen, and Wu}]{jia2019direct}
Ye~Jia, Ron~J Weiss, Fadi Biadsy, Wolfgang Macherey, Melvin Johnson, Zhifeng Chen, and Yonghui Wu. 2019.
\newblock Direct speech-to-speech translation with a sequence-to-sequence model.
\newblock \emph{arXiv preprint arXiv:1904.06037}.

\bibitem[{Kesiraju et~al.(2023{\natexlab{a}})Kesiraju, Bene{\v{s}}, Tikhonov, and {\v{C}}ernock{\'y}}]{kesiraju-etal-2023-systems}
Santosh Kesiraju, Karel Bene{\v{s}}, Maksim Tikhonov, and Jan {\v{C}}ernock{\'y}. 2023{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2023.iwslt-1.19} {{BUT} systems for {IWSLT} 2023 {M}arathi - {H}indi low resource speech translation task}.
\newblock In \emph{Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, pages 227--234, Toronto, Canada (in-person and online). Association for Computational Linguistics.

\bibitem[{Kesiraju et~al.(2023{\natexlab{b}})Kesiraju, Sarvaš, Pavlíček, Macaire, and Ciuba}]{kesiraju23_interspeech}
Santosh Kesiraju, Marek Sarvaš, Tomáš Pavlíček, Cécile Macaire, and Alejandro Ciuba. 2023{\natexlab{b}}.
\newblock \href {https://doi.org/10.21437/Interspeech.2023-2506} {Strategies for improving low resource speech to text translation relying on pre-trained asr models}.
\newblock In \emph{Interspeech 2023}, pages 2148--2152.

\bibitem[{Ko et~al.(2015)Ko, Peddinti, Povey, and Khudanpur}]{ko2015audio}
Tom Ko, Vijayaditya Peddinti, Daniel Povey, and Sanjeev Khudanpur. 2015.
\newblock Audio augmentation for speech recognition.
\newblock In \emph{Interspeech}, volume 2015, page 3586.

\bibitem[{Mbuya and Anastasopoulos(2023)}]{mbuya-anastasopoulos-2023-gmu}
Jonathan Mbuya and Antonios Anastasopoulos. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.iwslt-1.24} {{GMU} systems for the {IWSLT} 2023 dialect and low-resource speech translation tasks}.
\newblock In \emph{Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, pages 269--276, Toronto, Canada (in-person and online). Association for Computational Linguistics.

\bibitem[{Ortega et~al.(2024)Ortega, Zevallos, Ahmad, and Chen}]{ortega2024quespa}
John~E Ortega, Rodolfo~Joel Zevallos, Ibrahim~Sa’id Ahmad, and William Chen. 2024.
\newblock {QUESPA Submission for the IWSLT 2024 Dialectal and Low-resource Speech Translation Task}.
\newblock In \emph{Proceedings of the 21st International Conference on Spoken Language Translation (IWSLT 2024)}, pages 125--133.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and Zhu}]{papineni-etal-2002-bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock \href {https://doi.org/10.3115/1073083.1073135} {{B}leu: a method for automatic evaluation of machine translation}.
\newblock In \emph{Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics}, pages 311--318, Philadelphia, Pennsylvania, USA. Association for Computational Linguistics.

\bibitem[{Park et~al.(2019)Park, Chan, Zhang, Chiu, Zoph, Cubuk, and Le}]{park2019specaugment}
Daniel~S Park, William Chan, Yu~Zhang, Chung-Cheng Chiu, Barret Zoph, Ekin~D Cubuk, and Quoc~V Le. 2019.
\newblock Specaugment: A simple data augmentation method for automatic speech recognition.
\newblock \emph{arXiv preprint arXiv:1904.08779}.

\bibitem[{Popovi{\'c}(2017)}]{popovic-2017-chrf}
Maja Popovi{\'c}. 2017.
\newblock \href {https://doi.org/10.18653/v1/W17-4770} {chr{F}++: words helping character n-grams}.
\newblock In \emph{Proceedings of the Second Conference on Machine Translation}, pages 612--618, Copenhagen, Denmark. Association for Computational Linguistics.

\bibitem[{Post(2018)}]{post-2018-call}
Matt Post. 2018.
\newblock \href {https://doi.org/10.18653/v1/W18-6319} {A call for clarity in reporting {BLEU} scores}.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation: Research Papers}, pages 186--191, Brussels, Belgium. Association for Computational Linguistics.

\bibitem[{Post et~al.(2013)Post, Kumar, Lopez, Karakos, Callison-Burch, and Khudanpur}]{post2013improved}
Matt Post, Gaurav Kumar, Adam Lopez, Damianos Karakos, Chris Callison-Burch, and Sanjeev Khudanpur. 2013.
\newblock Improved speech-to-text translation with the fisher and callhome spanish-english speech translation corpus.
\newblock In \emph{Proceedings of the 10th international workshop on spoken language translation: papers}.

\bibitem[{Radhakrishnan et~al.(2023)Radhakrishnan, Agrawal, Prakash~Gohil, Praveen, Vinay~Dhopeshwarkar, and Pandey}]{radhakrishnan-etal-2023-sri}
Balaji Radhakrishnan, Saurabh Agrawal, Raj Prakash~Gohil, Kiran Praveen, Advait Vinay~Dhopeshwarkar, and Abhishek Pandey. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.iwslt-1.43} {{SRI}-{B}{'}s systems for {IWSLT} 2023 dialectal and low-resource track: {M}arathi-{H}indi speech translation}.
\newblock In \emph{Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, pages 449--454, Toronto, Canada (in-person and online). Association for Computational Linguistics.

\bibitem[{Robinson et~al.(2024)Robinson, Sun, Xiao, Bafna, Tan, Xu, Xinyuan, Kejriwal, Khudanpur, Murray et~al.}]{robinson2024jhu}
Nathaniel~Romney Robinson, Kaiser Sun, Cihan Xiao, Niyati Bafna, Weiting Tan, Haoran Xu, Henry~Li Xinyuan, Ankur Kejriwal, Sanjeev Khudanpur, Kenton Murray, and 1 others. 2024.
\newblock {JHU IWSLT 2024 Dialectal and Low-resource System Description}.
\newblock In \emph{Proceedings of the 21st International Conference on Spoken Language Translation (IWSLT 2024)}, pages 140--153.

\bibitem[{Shanbhogue et~al.(2023)Shanbhogue, Xue, Saha, Zhang, and Ganesan}]{vishnu-kudlu-shanbhogue-etal-2023-improving}
Akshaya Vishnu~Kudlu Shanbhogue, Ran Xue, Soumya Saha, Daniel Zhang, and Ashwinkumar Ganesan. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.iwslt-1.21} {Improving low resource speech translation with data augmentation and ensemble strategies}.
\newblock In \emph{Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, pages 241--250, Toronto, Canada (in-person and online). Association for Computational Linguistics.

\bibitem[{Wang et~al.(2020{\natexlab{a}})Wang, Tang, Ma, Wu, Okhonko, and Pino}]{wang-etal-2020-fairseq}
Changhan Wang, Yun Tang, Xutai Ma, Anne Wu, Dmytro Okhonko, and Juan Pino. 2020{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2020.aacl-demo.6} {Fairseq {S}2{T}: Fast speech-to-text modeling with fairseq}.
\newblock In \emph{Proceedings of the 1st Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics and the 10th International Joint Conference on Natural Language Processing: System Demonstrations}, pages 33--39, Suzhou, China. Association for Computational Linguistics.

\bibitem[{Wang et~al.(2020{\natexlab{b}})Wang, Tang, Ma, Wu, Popuri, Okhonko, and Pino}]{wang2020fairseq}
Changhan Wang, Yun Tang, Xutai Ma, Anne Wu, Sravya Popuri, Dmytro Okhonko, and Juan Pino. 2020{\natexlab{b}}.
\newblock Fairseq s2t: Fast speech-to-text modeling with fairseq.
\newblock \emph{arXiv preprint arXiv:2010.05171}.

\bibitem[{{Wikipedia contributors}(2025)}]{enwiki:1289695299}
{Wikipedia contributors}. 2025.
\newblock 2011 census of india --- {Wikipedia}{,} the free encyclopedia.
\newblock \url{https://en.wikipedia.org/w/index.php?title=2011_census_of_India&oldid=1289695299}.
\newblock [Online; accessed 1-June-2025].

\bibitem[{Zafar et~al.(2024)Zafar, Castaldo, Nayak, Haque, Gajakos, and Way}]{zafar2024setu}
Maria Zafar, Antonio Castaldo, Prashanth Nayak, Rejwanul Haque, Neha Gajakos, and Andy Way. 2024.
\newblock {The SETU-DCU submissions to IWSLT 2024 low-resource speech-to-text translation tasks}.
\newblock In \emph{Proceedings of the 21st International Conference on Spoken Language Translation (IWSLT 2024)}, pages 80--85.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[final]{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{graphicx}
\usepackage{booktabs}

\title{IIITH-BUT system for IWSLT 2025 low-resource Bhojpuri to Hindi speech translation}

\author{
  \textbf{Bhavana Akkiraju\textsuperscript{1}},
  \textbf{Aishwarya Pothula\textsuperscript{1}},
  \textbf{Santosh Kesiraju\textsuperscript{2}},
  \textbf{Anil Kumar Vuppala\textsuperscript{1}} \\
  \textsuperscript{1}International Institute of Information Technology, Hyderabad, India \\
  \textsuperscript{2}Speech@FIT, Brno University of Technology, Czechia \\
  \texttt{\{bhavana.akkiraju,aishwarya.pothula\}@research.iiit.ac.in} \\
  \texttt{kesiraju@fit.vut.cz,anil.vuppala@iiit.ac.in}
}

\begin{document}
\maketitle
\begin{abstract}
This paper presents the submission of IIITH-BUT to the IWSLT 2025 shared task on speech translation for the low-resource Bhojpuri-Hindi language pair. We explored the impact of hyperparameter optimisation and data augmentation techniques on the performance of the SeamlessM4T model fine-tuned for this specific task. We systematically investigated a range of hyperparameters including learning rate schedules, number of update steps, warm-up steps, label smoothing, and batch sizes; and report their effect on translation quality. To address data scarcity, we applied speed perturbation and SpecAugment and studied their effect on translation quality. We also examined the use of cross-lingual signal through joint training with Marathi and Bhojpuri speech data. Our experiments reveal that careful selection of hyperparameters and the application of simple yet effective augmentation techniques significantly improve performance in low-resource settings. We also analysed the translation hypotheses to understand various kinds of errors that impacted the translation quality in terms of BLEU.
\end{abstract}

\section{Introduction}

Speech translation (ST) transforms spoken language into written text in a different language, serving as a critical component in breaking down communication barriers. While significant advancements have been made for high-resource language pairs \cite{jia2019direct, bentivogli2021cascade}, developing effective ST systems for low-resource and dialectal languages remains challenging due to scarce parallel data, inconsistent orthography, and substantial linguistic variation.

We address the low-resource scenario of Bhojpuri speech to Hindi text translation within the Indian linguistic context. Despite being spoken by over 50 million people, Bhojpuri suffers from limited quality and diversity in available speech-text corpora \cite{enwiki:1289695299}. In contrast, Hindi possesses relatively abundant resources, making it an ideal target language for translation. Our system, developed for the IWSLT 2025 shared task, addresses these resource disparities through a combination of transfer learning from linguistically similar languages, data augmentation techniques, and hyperparameter optimization.

Recent end-to-end ST models, such as the Speech-to-Text Transformer \cite{wang2020fairseq} and SeamlessM4T \cite{barrault2023seamlessm4t} effectively replace traditional cascade pipelines by jointly modelling ASR and MT, reducing error propagation and latency. However, these approaches typically require substantial labelled data unavailable for most low-resource languages, often leading to overfitting and poor generalisation.

Our contributions include:
\begin{itemize}
    \item A systematic investigation of hyperparameter optimization for low-resource ST, identifying that configuration choices such as batch sizes, moderate label smoothing values, and extended warmup periods significantly impact performance on the Bhojpuri-Hindi language pair.
    \item An analysis of data augmentation techniques - SpecAugment \cite{park2019specaugment} and speed perturbation \cite{ko2015audio} for low-resource speech translation, demonstrating their effectiveness in expanding our training data by 3x and improving BLEU scores by an average of 2.1 points.
    \item An evaluation of cross-lingual transfer learning through joint fine-tuning with Marathi-Hindi data, empirically showing how linguistic similarities between related Indo-Aryan languages can be leveraged to improve low-resource speech translation performance.
\end{itemize}

The remaining  of this paper is organized as follows: Section~\ref{sec:lit} provides a comprehensive review of related work, Section~\ref{sec:model} describes our system including hyperparameter optimization, data augmentation techniques, and joint fine-tuning approach, Section~\ref{sec:exp} details our experimental setup, Section~\ref{sec:results} presents results and analysis, and Section~\ref{sec:conc} concludes with future directions.

\section{Related work}
\label{sec:lit}
Low-resource speech translation (ST) has garnered significant attention through IWSLT shared tasks \cite{agrawal-etal-2023-findings,ahmad-etal-2024-findings,gow-smith-etal-2023-naver}. The field has evolved from traditional pipeline approaches \cite{post2013improved}  to end-to-end architectures such as Listen Attend and Spell (LAS) \cite{berard2016listen}, fairseq S2T \cite{wang-etal-2020-fairseq,e-ortega-etal-2023-quespa},\cite{radhakrishnan-etal-2023-sri}, and transfer learning based methods \cite{kesiraju23_interspeech,kesiraju-etal-2023-systems}. In \cite{mbuya-anastasopoulos-2023-gmu}, explored fine-tuning self-supervised models by incorporating a linear layer for the ST task, which streamlined workflows while maintaining specialized strategies for low-resource scenarios. \cite{vishnu-kudlu-shanbhogue-etal-2023-improving}, implemented various data augmentation techniques including audio stretching, back-translation, and paraphrasing.

Contemporary approaches to low-resource ST can be categorized into several methodological frameworks. The SETU-DCU submission \cite{zafar2024setu} enhanced ST robustness through CTC loss integration and rigorous data cleaning protocols. \cite{post2013improved} incorporated pseudo-labelling techniques to expand their training corpus. The JHU IWSLT 2024 system \cite{robinson2024jhu} demonstrated the efficacy of Whisper-style large models with domain-adaptive pretraining methodologies. Meanwhile, the QUESPA team \cite{ortega2024quespa} implemented ensemble decoding with cross-lingual knowledge transfer mechanisms. SeamlessM4T \cite{barrault2023seamlessm4t}  presents a comprehensive approach handling ASR, MT, and ST across more than 100 languages, though its performance on genuinely low-resource languages necessitates substantial adaptation strategies.

Our approach focuses on systematically exploring hyperparameter optimization and data augmentation techniques for low-resource speech translation. Unlike previous work that often applies general strategies such as default hyperparameter values, generic data augmentation, standard transfer learning without language-specific considerations and using default model architectures and training strategies without adaptation to low-resource constraints, we conduct a comprehensive investigation specifically tailored to the challenges of the Bhojpuri-Hindi language pair. Our experimentation includes a detailed analysis of learning rates, batch sizes, label smoothing values, and warmup periods, and  used two data augmentation techniques (Speed Perturb, SpecAug). Additionally, we examine how cross-lingual transfer from Marathi can supplement these optimizations, leveraging the linguistic proximity between these Indo-Aryan languages.

\section{Methodology}
\label{sec:model}
\subsection{Model architecture}
Our experiments used SeamlessM4T as the backbone. We experimented with medium (1.2B parameters) and large (2.3 B parameters) variants. The medium consists of a 24-layer conformer speech encoder, 12-layer Transformer text decoder, with 1024-dimensional hidden states, and 16 attention heads. 

\subsection{Fine-tuning}
For most of the experiments, we fine-tuned all the parameters on the target language pair, i.e. Bhojpuri--Hindi. We also conducted experiments where fine-tuned on both language pairs Marathi--Hindi and Bhojpuri--Hindi. This model was further fine-tuned for few epochs for the target pair Bhojpuri--Hindi.

\subsection{Evaluation}
We used the standard objective metrics BLEU \cite{papineni-etal-2002-bleu} and chrF++ \cite{popovic-2017-chrf}  implemented in sacrebleu \cite{post-2018-call}  to objectively evaluate the translation quality against the reference. 

\section{Experiments}
\label{sec:exp}

\subsection{Datasets}
We used only the official IWSLT 2025 shared task dataset for Bhojpuri → Hindi speech translation:
\begin{table}[ht]
\centering
\begin{tabular}{lrr}
\toprule
\textbf{Dataset} & \textbf{Duration (hrs)} & \textbf{No. of utterances} \\
\midrule
Training  & 20.00 & 10,171 \\
Dev & 2.07 & 1,056 \\
Test & 0.87 & 750 \\
\bottomrule
\end{tabular}
\caption{Data split statistics for Bhojpuri-Hindi ST task.}
\label{tab:data_stats}
\end{table}
For our multilingual experiments, we incorporated an additional Marathi $\rightarrow$ Hindi\footnote{\href{https://github.com/panlingua/iwslt2025_mr-hi}{IWSLT Marathi-Hindi Dataset}} parallel corpus from IWSLT, consisting of 16 hours (7,990 utterances) of training data, 3.6 hours (2,103 utterances) of development data, and 0.45 hours (286 utterances) of test data.

\subsection{Implementation}
Our code was based on the original seamless library written in PyTorch.
Our experiments used two NVIDIA A100 GPUs (40GB each), employing data parallel training, synchronised batch normalisation, and FP16 mixed precision for optimal computational efficiency.

\subsection{Pre-trained model selection}
Here we report the fine-tuning results for Seamless medium and large v2 variants. We observed that the medium model consistently yielded better translation results than the large one. Hence, we used the medium variant for all subsequent experiments and analysis.

\begin{table}[!ht]
    \centering
    \begin{tabular}{lrrr}
    \toprule
    \textbf{Model} & lr & BLEU & chrF++ \\ \midrule
     Medium & 1e-5 & \textbf{30.5} & \textbf{54.6} \\
     Large  & 1e-5  &25.5 & 50.5\\
     
     \bottomrule
    \end{tabular}
    \caption{Translation scores on dev set after fine-tuning seamless medium and large variants on  Bhojpuri--Hindi.}
    \label{tab:med_vs_large}
\end{table}

\subsection{Hyperparameter optimisation}
Our hyperparameter optimisation investigation focused on several critical training parameters. The \textbf{learning rate} (LR) was evaluated across three settings (1e-6, 1e-5, and 2e-5), with the moderate rate of \textbf{1e-5} consistently yielding the best performance as shown in Table~\ref{tab:lr_comp}, achieving a balance between training stability and domain adaptation. For \textbf{label smoothing} (LS), we tested values of 0, 0.1, and 0.2, with \textbf{0.1} offering the best generalisation and robustness (Table~\ref{tab:ls_comp}). We explored \textbf{warmup steps} (100, 250, 350, and 400) to stabilise the learning rate schedule, finding \textbf{250 steps} produced optimal convergence without extending training time. To mitigate overfitting, we experimented with \textbf{early stopping patience} values (5, 10, and 20 epochs), where \textbf{10 epochs} struck the best trade-off between overtraining and early termination. Finally, we examined \textbf{training batch size} (5, 10, 32, and 64), with a batch size of \textbf{32} demonstrating the most favourable performance (Table~\ref{tab:lr_comp}), balancing gradient update stability with computational efficiency.
\begin{table}[th]
  \centering
  \begin{tabular}{lrrr}
    \toprule
    \textbf{LR} & \textbf{batch size} & \textbf{BLEU} & \textbf{chrF++} \\
    \midrule
    1e-5 & 5   & 30.5 & 54.5 \\
    1e-5 & 10  & 33.8 & 56.9 \\
    1e-5 & 32 & \textbf{35.1} & 58.2 \\ 
    \hline
    1e-6 & 5   & 18.2 & 48.4 \\
    1e-6 & 10  & 20.1 & 49.1 \\
    1e-6 & 32  & 26.7 & 51.2    \\
    \bottomrule
  \end{tabular}
  \caption{Effect of learning rate (LR) and training batch size (batch size) on Bhojpuri-Hindi fine-tuning performance}
  \label{tab:lr_comp}
\end{table}

\begin{table}[th]
  \centering
  \begin{tabular}{lrr}
    \toprule
    \textbf{LS} & \textbf{BLEU} & \textbf{chrF++} \\
    \midrule
    0.0    & 30.9         & 55.3 \\
    0.1  & \textbf{33.8}  & \textbf{56.9} \\
    0.2  & 31.8        & 56.6 \\
    \bottomrule
  \end{tabular}
  \caption{Effect of label smoothing (LS) on Bhojpuri--Hindi fine-tuning performance with LR = 1e-5 and batch size = 10.}
  \label{tab:ls_comp}
\end{table}

\subsection{Data augmentation}
To address the limited training data for the Bhojpuri-Hindi language pair, we implemented two established speech augmentation techniques.

\subsubsection{SpecAugment}
We applied spectrogram masking with time masks (max 30 frames) and frequency masks (max 30 mel-frequency bins), creating diverse variations that forced the model to rely on broader contextual information rather than specific acoustic features.

\subsubsection{Speed perturbation}
We implemented speed factors of 0.9x, 1.0x, and 1.1x to simulate different speaking rates without changing pitch, effectively tripling our training data with realistic variations and improving the robustness of the model to natural speaking rate differences among Bhojpuri speakers.

By combining these complementary methods, we expanded the diversity of training data without requiring additional recordings. The impact on translation quality is shown in Table~\ref{tab:sp}.

\begin{table}[th]
  \centering
  \begin{tabular}{ccr}
    \toprule
    \textbf{SP} & \textbf{SA} & \textbf{BLEU}  \\
    \midrule
    False & False & 31.8 \\
    \textbf{False} & \textbf{True} & \textbf{33.7} \\
    True & False &  32.7 \\
    True & True & 32.4  \\
    \bottomrule
  \end{tabular}
  \caption{Effect of Speed perturb (SP) and  SpecAugment (SA) on Bhojpuri--Hindi fine-tuning performance with LR = 1e-5, batch size = 10, and LS = 0.1}
  \label{tab:sp}
\end{table}


\subsection{Joint-finetuning approach}
We implemented cross-lingual transfer learning by integrating Marathi-Hindi and Bhojpuri-Hindi language pairs into a unified training framework. This approach leverages the linguistic similarities between these Indo-Aryan languages, which share phonological characteristics, lexical resources, and syntactic structures.

Our methodology combined the Marathi-Hindi parallel corpus with the limited Bhojpuri-Hindi dataset during SeamlessM4T model adaptation, thereby expanding the training data while introducing linguistically relevant patterns from Marathi. To mitigate catastrophic forgetting, we implemented sequential fine-tuning with an initial joint training phase followed by Bhojpuri-only fine-tuning. We tested this approach under three conditions: one epoch, two epochs, and training until convergence, as shown in Table~\ref{tab:jointft}.

\begin{table*}[th]
\centering
\begin{tabular}{lrrrrrrrr}
\toprule
\textbf{LR} & \textbf{LS} & \textbf{Batch size} & \textbf{SP} & \textbf{SA} & \textbf{Warm up steps} & \textbf{Patience} & \textbf{Beam size} & \textbf{BLEU} \\
\midrule
1e\_5 & 0.1 & 10 & False & False & 100 & 5 & 5 & 33.1 \\
\midrule
1e\_5 & 0.1 & 10 & False & False & 100 & 5 & 10 & 33.8 \\
1e\_5 & 0.1 & 32 & False & False & 100 & 5 & 10 & 34.0 \\
\midrule
1e\_5 & 0.1 & 32 & False & True & 250 & 5 & 10 & 35.3 \\
\textbf{1e\_5} & \textbf{0.1} & \textbf{32} & \textbf{False} & \textbf{True} & \textbf{250} & \textbf{10} & \textbf{10} & \textbf{36.4} \\
1e\_5 & 0.1 & 32 & False & True & 250 & 20 & 10 & 35.6 \\
\bottomrule
\end{tabular}
\caption{BLEU scores for various hyperparameter configurations during fine-tuning of SeamlessM4T. We varied the learning rate (LR), label smoothing (LS), batch size, speed perturbation (SP), SpecAugment (SA), warm-up steps, early stopping patience, and beam size. The highest BLEU score (36.41) was obtained with SA enabled, patience set to 10, and a beam size of 10.}
\label{tab:final}
\end{table*}

\begin{table}[th]
\centering
\begin{tabular}{lrr}
\toprule
\textbf{Strategy} & \textbf{Epochs} & \textbf{BLEU}  \\
\midrule
Joint finetuning (JF) & Convergence & 34.6 \\
\midrule
 \textbf{JF + monolingual bhoj} &\textbf{1} &  \textbf{36.0} \\
 JF + monolingual bhoj & 2 & 35.6  \\
 JF + monolingual bhoj & Convergence & 35.4 \\
\bottomrule
\end{tabular}
\caption{BLEU scores for contrastive model with the same configuration as highest BLEU in Table~\ref{tab:final} }
\label{tab:jointft}
\end{table}

\section{Results and analysis}
\label{sec:results}
This section presents an analysis of the experimental results obtained from our primary and contrastive models. The \textbf{primary model} encompasses an optimized combination of hyperparameter configuration and data augmentation techniques that yielded the highest BLEU score in our evaluations.

As demonstrated in Table~\ref{tab:sp}, the application of SpecAugment during fine-tuning produced superior performance compared to other augmentation strategies. During inference, we systematically evaluated translation quality across multiple beam search configurations (sizes 1, 5, and 10) to determine the optimal decoding approach.

Table~\ref{tab:final} represents a comprehensive comparison of various experimental configurations and their corresponding performance metrics. Increasing the beam size from 5 to 10 while maintaining all other parameters constant yielded a modest improvement in translation quality. Subsequently, with beam size fixed at 10, enlarging the training batch size from 10 to 32 further enhanced performance by 0.21 BLEU points. The combination of beam size 10, batch size 32, increased warmup steps from 100 to 250, and the introduction of SpecAugment collectively improved the BLEU score from 34.01 to 35.38. Our optimal configuration, which additionally increased early stopping patience from 5 to 10, achieved the highest performance with 36.41 BLEU. Notably, further extending patience to 20 epochs resulted in performance degradation, suggesting potential overfitting.

For our \textbf{contrastive model}, we implemented a multilingual fine-tuning strategy that jointly trained on Marathi and Bhojpuri data using the hyperparameter configuration that previously achieved the highest performance (36.4 BLEU). This multilingual model initially underperformed compared to our monolingual system, potentially due to catastrophic forgetting. To mitigate this issue, we conducted additional fine-tuning on Bhojpuri data exclusively for various epoch counts. 

As shown in Table~\ref{tab:jointft}, performance peaked after a single epoch of Bhojpuri-specific fine-tuning and subsequently declined with additional epochs, suggesting that extended training on previously observed data may lead to overfitting. Consequently, our contrastive submission consisted of the joint Marathi-Bhojpuri model with one additional epoch of Bhojpuri--Hindi exclusive fine-tuning, which produced results comparable to our optimised monolingual configuration.

Table~\ref{tab:testresult} presents the BLEU scores obtained on both test and development datasets. In the IWSLT 2025 shared task \cite{iwslt2025}, the highest reported BLEU score was 10.7, representing a significant decrease compared to IWSLT 2024 \cite{ahmad-etal-2024-findings}, where scores reached approximately 24.4. Our primary and contrastive models achieved BLEU scores of 9.9 and 10.2 respectively on the test set, while demonstrating substantially higher performance on the development set with scores of 36.4 and 36.0. The considerable performance gap between development and test sets suggests potential domain mismatch between the datasets or possible data quality issues in the test set, warranting further investigation.

\begin{table}[th]
\centering
\begin{tabular}{lrr}
\toprule
\textbf{Model} & \textbf{Dev BLEU} & \textbf{Test BLEU}  \\
\midrule
Primary   & 36.4 & 9.9   \\
\midrule
Contrastive & 36.0 &  10.2 \\
\bottomrule
\end{tabular}
\caption{BLEU scores for primary and contrastive models using the same configuration as highest BLEU scores in Table~\ref{tab:final} and Table~\ref{tab:jointft} for both dev and test dataset }
\label{tab:testresult}
\end{table}

\subsection{Error analysis}
Our systematic examination of translation outputs revealed several factors in the development dataset that affected ST performance. Analysis of audio-transcript relationships identified multiple inconsistencies impacting model performance. We observed three key patterns when comparing reference and target word counts: (1) When reference counts exceeded target counts, low BLEU scores often resulted from incomplete audio recordings paired with complete reference transcripts, audio-transcript misalignment, or redundant reference content; (2) Equal word counts with low BLEU scores frequently corresponded with noisy recordings; (3) Cases where target counts exceeded reference counts typically involved recordings with significant acoustic interference.

Numerical content presented particular challenges. We identified inconsistent representation formats (e.g., "8 crores 74 lakhs" in audio versus "87.4 lakhs" in text), incomplete numerical transcription (e.g., in audio, numbers are spoken in English as "Fifteen" whereas in reference text they appear in Hindi as "Pandrah(hindi)"), and instances where equal numerical representation corresponded with degraded audio quality. These findings highlight the importance of audio-transcript alignment and standardized numerical representation in speech translation datasets, particularly for low-resource language evaluation.

\section{Conclusion}
\label{sec:conc}
Our submission to the IWSLT 2025 evaluation campaign for low-resource and dialectal speech translation advances Bhojpuri--Hindi ST through a combination of hyperparameter optimisation, data augmentation, and cross-lingual joint fine-tuning. By leveraging the SeamlessM4T medium model (1.2B parameters) and systematically exploring optimal training configurations, we demonstrate significant performance gains despite the challenges posed by limited parallel data.
Our results show that even established techniques like SpecAugment and speed perturbation, when carefully implemented, can lead to substantial improvements in low-resource speech translation tasks, expanding our effective training data threefold. Additionally, we found that joint training with Marathi—a linguistically related Indo-Aryan language—followed by sequential Bhojpuri-specific adaptation provides an effective strategy to mitigate data sparsity and improve generalisation across diverse speech patterns.

For future work, we plan to explore more sophisticated augmentation techniques such as noise injection, pitch shifting, and cross-speaker synthesis. We also intend to investigate self-supervised pretraining on monolingual Bhojpuri speech data and further extend our cross-lingual approach to additional Indo-Aryan languages. As low-resource ST continues to evolve, we believe that modular, linguistically informed adaptation pipelines will play a key role in advancing the real-world applicability of such systems for under represented language communities.

\section{Acknowledgements}

Santosh Kesiraju was supported by Ministry of Education, Youth and Sports of the Czech Republic (MoE) through the OP JAK project ``Linguistics, Artificial Intelligence and Language and Speech Technologies: from Research to Applications'' (ID:CZ.02.01.01/00/23\_020/0008518).

\bibliography{anthology,custom}
\end{document}